OpenStudy (anonymous):

which expression is equal to the expression below ? (6 *9) 5 a. 15 to the 5th b. 5* (6*9) c. 6 to the 5th + 9 to the 5th d. 6 to the 5th times 9 to the 5th ?

OpenStudy (ash2326):

We have \[(6\times 9) 5\] The associative property of multiplication states that when multiplying three or more real numbers, the product is always the same regardless of their grouping. That is, \[(a × b) × c = a × (b × c)\] so \[(6\times 9) 5=6( 9 \times 5)\]

OpenStudy (hoblos):

i think "to the 5th" means power 5 not times 5 !!

OpenStudy (ash2326):

Sorry I made an error \[(6\times 9) 5=>>>5( 6\times 9) \]

OpenStudy (ash2326):

@blanca is it \(\huge (6\times 9)^5\) ?

OpenStudy (hoblos):

(6 *9)^5 = (6^5)(9^5) d